OpenStudy (anonymous):

What is the solution to the equation log base 3 of x – log base 3 of 2 = 2 ? a)x = 18 b)x = 12 c)x = 9/2 d)x = 4/3

OpenStudy (anonymous):

start with \[\log_3(\frac{x}{2})=2\] tjem \[\frac{x}{2}=3^2\] and go from there